# On De Giorgi’s Conjecture of Nonlocal approximations for free-discontinuity problems: The symmetric gradient case

### v1

@@ Abstract.
We prove that E. De Giorgi’s conjecture for the nonlocal approximation of free-discontinuity problems extends to the case of functionals defined in terms of the symmetric gradient of the admissible field. After introducing a suitable class of continuous finite-difference approximants, we show the compactness of deformations with equibounded energies, as well as their Gamma-convergence. The compactness analysis relies on a generalization of a Fréchet-Kolmogorov approach previously introduced by two of the authors. An essential difficulty is the identification of the limiting space of admissible deformations. We show that if the approximants involve superlinear contributions, a limiting GSBD representation can be ensured, whereas a further integral geometric term appears in the limiting functional in the linear case. We eventually discuss the connection between this latter setting and an open problem in the theory of integral geometric measures.
@@ 1. Introduction
Free-discontinuity problems and their approximation by means of Sobolev formulations, discrete descriptions, or nonlocal functionals are a thriving research area owing to their broad scope of applications, ranging from image reconstruction, to the modeling of failure phenomena in continuum mechanics. A milestone in this direction was a conjecture formulated by E. De Giorgi, and dealing with the approximation via Gamma-convergence of the Mumford-Shah functional by a sequence of nonlocal counterparts. Starting from the seminal proof of this result by M. Gobbino in [26], such analysis has paved the way for further generalizations in [17, 27], eventually leading from [25] to variational studies on point clouds (see, e.g., [6, 11, 19]), as well as to machine-learning applications [10]. All the aforementioned contributions deal with free-discontinuity problems involving the full distributional gradient of the admissible maps. A question that, to the authors’ knowledge, was so far left open, was whether similar nonlocal approximation techniques would also prove successful for the study of free-boundary problems involving more general differential operators.
In this paper we show that an analogue approximation result to that conjectured by E. De Giorgi and proved by M. Gobbino holds in the case of free-boundary problems defined in terms of the symmetric gradient of the admissible deformations. To be precise, we introduce a new class of continuous finite-difference approximations of linearized Griffith functionals, and we rigorously derive their $\Gamma$-convergence as well as the compactness of their minimizing sequences.
In order to describe our findings, we first recall the results [26], where the sequence of functionals
@@
In (1.2), the acronym ${\rm GSBV}$ stands for the space of functions with generalised special bounded variation (roughly speaking, maps whose truncations are functions of bounded variations with distributional gradients exhibiting null Cantor part), cf. [4]. Note that the multiplicative constants in (1.2) depend only on the choice of the integrand $\arctan(x)$ and of the weight function $e^{-|x|^{2}}$.
@@
The structure of (1.3) aligns with the principles of linearized theories in continuum mechanics, where only the symmetric part of the gradient contributes to the deformation energy. This consideration naturally motivates the assumption that only the component of $u(x^{\prime})-u(x)$ projected along the direction $x-x^{\prime}$ should provide energy contributions. By setting $\varepsilon\xi:=x^{\prime}-x$, when the differences $(u(x+\varepsilon\xi)-u(x))\cdot\xi$ are relatively small, the functionals in (1.3) behave like pure bulk energies. For large values of $(u(x+\varepsilon\xi)-u(x))\cdot\xi$, instead, the energies in (1.3) saturate, effectively detecting and penalizing the size of the discontinuities of $u$.
An important remark concerns the fact that the energies in (1.1) only approximate the Mumford-Shah functional under suitable additional assumptions on the geometry of $\Omega$, such as Lipschitz regularity (see [26, Remark 7.1]). Such requirements may not be fully consistent when handling free-boundary problems where discontinuities are already present within the material. Consider, for example, the case of a crack-initiated domain, where the presence of an $(n-1)$-dimensional set $\Gamma\subset\Omega$, representing an initial crack makes the set $\Omega\setminus\Gamma$ not Lipschitz.
@@
In the expression above, $\mathscr{B}$ represents the class of all finite families of pairwise disjoint open balls contained in $\Omega$. This supremum procedure over $\mathscr{B}$ allows the functional $\mathcal{F}^{p}_{\varepsilon}$ to bypass irregularities within the domain $\Omega$. In fact, by choosing $p=1$, the strict superaddiditity of the set function $F_{\varepsilon,\xi}(u,\cdot)$ leads to the strict inequality $\mathcal{F}^{1}_{\varepsilon}<\mathcal{F}_{\varepsilon}$. For $p>1$, the presence of the $L^{p}$-norm in the definition (1.4) should be viewed as a regularizing effect, and its precise role will be explained in more detail below.
Our main results are provided by the following two theorems.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be open, let $p>1$, and let $\{u_{\varepsilon}\}_{\varepsilon>0}\subset L^{0}(\Omega;\mathbb{R}^{n})$ be such that
@@
$\liminf_{k\to\infty}\mathcal{F}^{p}_{\varepsilon_{k}}(u_{k},\Omega)\geq\int_{\Omega}\varphi_{p}(e(u))\textnormal{d}x+\beta_{p}\mathcal{H}^{n-1}(J_{u}\cup\partial^{*}A),$ (1.7)
@@
where $\varphi_{p}\colon\mathbb{M}^{n\times n}_{sym}\to[0,\infty)$ is $2$-homogeneous and $\beta_{p}$ is a positive constant.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be open. For every $p>1$ the family of functionals $\{\mathcal{F}^{p}_{\varepsilon}\}_{\varepsilon>0}$ $\Gamma$-converges in $L^{0}(\Omega;\mathbb{R}^{n})$ to the Griffith-type functional $\mathcal{F}^{p}$ defined as
@@
In proving Theorems 1.1 and 1.2 we addressed two main hurdles. The first challenge, common to both theorems, involves the characterization of the domain of the $\Gamma$-limit. The second difficulty concerns the compactness result in Theorem 1.1, where no integrability assumptions are imposed on the family $\{u_{\varepsilon}\}_{\varepsilon}$.
The assumption $p>1$ plays a key role in tackling the first problem. Broadly speaking, the non-local nature of the approximating functional $\mathcal{F}^{p}_{\varepsilon}$ results in a limiting function space consisting of measurable vector fields that exhibit generalized bounded deformation in a weaker sense. To explain this phenomenon more precisely, recall that $u\in{\rm GBD}(\Omega)$ if and only if $u$ is a measurable vector field and there exists a finite Radon measure $\lambda$ on $\Omega$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the generalized directional variation $\hat{\mu}_{u}^{\xi}$ of $u\cdot\xi$, as introduced in [21, Definition 4.10], satisfies:
By a slicing argument, a similar approach to that proposed by M. Gobbino in [26] gives that a measurable vector field $u\colon\Omega\to\mathbb{R}^{n}$ belongs to the domain of $\mathcal{F}^{1}$ if and only if the condition in (1.9) holds in a much weaker form, namely only after averaging with respect to $\xi$,
The fact that (1.9) is replaced in our setting by (1.10) is a huge obstacle to establish the structural properties of the limiting function space. In particular, the technique presented in [21] for deriving the GBD-structural properties fails, and we need to rely on the more recent integral geometric approach introduced in [2]. The main difference from the approach in [2] lies in the role of the flatness of $\mathbb{R}^{n}$, where rescaling and translating the domain preserve straight lines as geodesics, maintaining invariance in the associated GBD-space. This property does not generally hold on a Riemannian manifold. In Proposition 3.13, we leverage this feature to avoid relying on a Korn-Poincaré-type inequality to estimate the dimensionality of the set on which the jump of all one-dimensional slices of $u$ concentrate. In this direction, the exponent $p$ in $(1,\infty)$ in (1.4) acts as a regularization, for it provides in the limit an interpolation between the two conditions in (1.9) and (1.10). The assumption that $p>1$ in Theorems 3.6 and 3.7 is thus instrumental to infer essential structural properties for the maps in the domain of the limiting energy, particularly those related to the slicing of the jump set. Ultimately, we can conclude in Theorem 3.1 that the domain of $\mathcal{F}^{p}$ coincides precisely with ${\rm GSBD}(\Omega)$. For the detailed definitions of the relevant quantities and the proof of the structural properties, we refer the reader to Sections 2 and 3, respectively. It is important to observe that the significant non-linearity in the definition of $\hat{\mu}^{\xi}_{u}$ prevents us from controlling the averaged measure in (1.10) through finitely many $\hat{\mu}^{\xi_{i}}_{u}$ for $i=1,\dotsc,N$, regardless of the chosen finite set of directions $\{\xi_{1},\dotsc,\xi_{N}\}$. Such a bound can only be established after deriving the structural properties of the limiting maps. This contrasts sharply with the GSBV setting analyzed in [26, Lemma 4.1]: in that case, although truncations were also employed to define the space, its structure remained deeply rooted in the theory of distributions.
To clarify why the case $p=1$ remains unresolved, denote by ${\rm GBV}^{\mathcal{E}}(\mathbb{R}^{n})$ the space of measurable vector fields $u\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ such that, for $\mathcal{H}^{n-1}$-almost every $\xi\in\mathbb{S}^{n-1}$ and for $\mathcal{H}^{n-1}$-almost every $y\in\Pi^{\xi}$ (where $\Pi^{\xi}$ is the orthogonal hyperplane to $\xi$ passing through the origin), the one-dimensional slice $\hat{u}^{\xi}_{y}$ belongs to $BV_{\mathrm{loc}}(\mathbb{R})$, together with the condition in (1.10). For precise definitions of the relevant quantities, we refer to Section 2. For every Borel set $B\in\mathbb{R}^{n}$ and every $y\in\mathbb{R}^{n}$, denote further $B^{\xi}_{y}:=\{t\in\mathbb{R}:\,y+t\xi\in B\}.$
The setting $p=1$ is related to the following question.
Open Problem.
Let $u\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ be a measurable vector field, and define the Borel regular measure $\mathscr{I}_{u,1}$ on $\mathbb{R}^{n}$ as
$\mathscr{I}_{u,1}(B):=\int_{\mathbb{S}^{n-1}}\bigg{(}\int_{\Pi^{\xi}}\sum_{t\in B^{\xi}_{y}}|[\hat{u}^{\xi}_{y}](t)|\wedge 1\,\mathrm{d}\mathcal{H}^{n-1}(y)\bigg{)}\mathrm{d}\mathcal{H}^{n-1}(\xi),\quad B\subset\mathbb{R}^{n}\text{ Borel}.$
Is it true that if $u\in{\rm GBV}^{\mathcal{E}}(\mathbb{R}^{n})$, then $\mathscr{I}_{u,1}$ is an $(n-1)$-rectifiable measure?
This problem is quite delicate, as it hinges on the ability to deduce the rectifiability of a measure expressed in an integral geometric form. As P. Mattila showed in [29], for some explicit counterexamples the answer is a negative one. The question whether similar constructions would apply to the measure described above currently remains open. For a more detailed discussion on this topic, we refer the interested reader to Section 6. Without knowing the answer to the open problem raised above, our $\Gamma$-convergence result for $p=1$ reads as follows.
Theorem 1.3 ($\Gamma$-convergence for $p=1$).
Let $\Omega\subset\mathbb{R}^{n}$ be open. The family of functionals $\{\mathcal{F}^{1}_{\varepsilon}\}_{\varepsilon}$ $\Gamma$-converges in $L^{0}(\Omega;\mathbb{R}^{n})$ to the Griffith-type functional $\mathcal{F}^{1}$ defined as
$\mathcal{F}^{1}(u,\Omega):=\begin{cases}\displaystyle\int_{\Omega}\varphi(e(u))\textnormal{d}x+\beta\mathcal{H}^{n-1}(J_{u})+\mathcal{I}_{u}^{s}(\Omega)&\text{for $u\in{\rm GSBV^{\mathcal{E}}}(\Omega)$,}\\[5.69054pt] \displaystyle\infty&\text{otherwise in $L^{0}(\Omega;\mathbb{R}^{n})$,}\end{cases},$
where $\varphi\colon\mathbb{M}^{n\times n}_{sym}\to[0,\infty)$ is $2$-homogeneous and $\beta$ is a positive constant, while the measure $\mathcal{I}_{u}^{s}$ is absolutely continuous with respect to the purely $(n-1)$-unrectifiable part of $\mathscr{I}_{u,1}$, and, in particular, gives zero mass to every $\sigma$-finite set with respect to $\mathcal{H}^{n-1}$.
Once the identification of the domain of the $\Gamma$-limit was obtained, the second main issue which we had to face was the compactness result stated in Theorem 1.1. In this regard, we recall some related works  [23, 28, 32], inspired by the corresponding results on the Mumford-Shah functional [7, 30], which provide a nonlocal approximation of a class of Griffith-type functionals in terms of a bulk energy of the form
featuring the convolution of a volume density $W(e(u))$ for a Sobolev map $u$ and a potential $f$ whose behavior is similar to that of $\arctan$ in (1.3). We further mention [1, 18], which instead proposed a discrete finite-difference approximation in the spirit of [5, 13, 14, 31], with a finite range of interaction.
Comparing our setting with those in the above literature, the main difficulty to overcome in the study of (1.3) was the lack of information on the gradient (or symmetric gradient) of $u$, which is, a priori, only a measurable function. This was a major obstacle to establish the compactness result in Theorem 1.1. Indeed, we could not directly apply the well-known compactness results of [15, 16] nor modify their technique, which builds upon a slicing argument and the existence of a generalised symmetric gradient/generalised variation. Surprisingly enough, we could still rely on a Fréchet-Kolmogorov approach, which amounts to prove equi-continuity of translations. This technical step is provided in the proof of Theorem 1.1 and shares common ideas with [3], where the proof of the compactness theorem of [15] was first revisited, avoiding the use of Korn and Korn-Poincaré inequalities. Once again, the lack of a symmetric gradient and the structure of $\mathcal{F}_{\varepsilon}$, which features an extra integration over the directions $\xi\in\mathbb{R}^{n}$, do not allow us to select a preferred basis to control translations, as it was the case in [3]. Nevertheless, the slicing properties of the functionals $\mathcal{F}_{\varepsilon}$ make it possible to control the translations of the maps $\arctan(u_{\varepsilon}(x)\cdot\xi)$ both with respect to $x$ and $\xi$. Compactness is thus recovered by Fréchet-Kolmogorov Theorem. Eventually, combining Theorem 1.1 with the $\Gamma$-convergence of Theorem 1.2, we further deduce convergence of minimisers of $\mathcal{F}_{\varepsilon}^{p}$ to minimisers of $\mathcal{F}^{p}$ under suitable Dirichlet boundary conditions (cf. Theorem 5.2).
@@
Organization of the paper. The paper is organized as follows. In Section 2, we collect a few preliminary definitions and results. Section 3 is devoted to establish the main properties of the spaces of limiting deformations. Sections 4 and 5 are devoted to the proofs of our compactness, and $\Gamma$-convergence result, respectively. Eventually, the case $p=1$ is the subject of Section 6.
@@
Outlook. The analysis provided in this work offers the opportunity to explore several further research lines. On the one hand, the slicing technique has been successfully employed in other finite difference approximations, such as [8, 12, 25] for the Total Variation, the Mumford-Shah, and the perimeter functionals. On the other hand, the use of a slicing argument for the study of the variational limit restricts the choice of the approximating sequence $\mathcal{F}_{\varepsilon}$, which in turn determines the class of admissible densities $\varphi_{p}$. The study of more general approximations and of integral representation formulas in the spirit of [17] will be the subject of future investigation.

### v2

@@ Abstract.
We prove that E. De Giorgi’s conjecture for the nonlocal approximation of free-discontinuity problems extends to the case of functionals defined in terms of the symmetric gradient of the admissible field. After introducing a suitable class of continuous finite-difference approximants, we show the compactness of deformations with equibounded energies, as well as their Gamma-convergence. The compactness analysis is a crucial hurdle, which we overcome by generalizing a Fréchet-Kolmogorov approach previously introduced by two of the authors. A second essential difficulty is the identification of the limiting space of admissible deformations, since a control on the directional variations is, a priori, only available in average. A limiting representation in GSBD is eventually established via a novel characterization of this space.
@@ 1. Introduction
Free-discontinuity problems and their approximation by means of Sobolev formulations, discrete descriptions, or nonlocal functionals are a thriving research area owing to their broad scope of applications, ranging from image reconstruction, to the modeling of failure phenomena in continuum mechanics. A milestone in this direction was a conjecture formulated by E. De Giorgi and proven by M. Gobbino [31], dealing with the approximation via $\Gamma$-convergence of the Mumford-Shah functional by a sequence of nonlocal counterparts. Such analysis has then paved the way for further generalizations in [22, 32]. All the aforementioned contributions deal with free-discontinuity problems involving the full distributional gradient of the admissible maps. A question that, to the authors’ knowledge, was so far left open, was whether similar nonlocal approximation techniques would also prove successful for the study of free-boundary problems involving more general differential operators.
In this paper we initiate the study of continuous finite-difference approximations of linearized Griffith functionals. Aside from purely mathematical interest, such inquiries are deeply rooted in the recent research lines which are developing in image processing and data science. Motivated by applications in Magnetic Resonance Imaging (MRI) or Positron Emission Tomography (PET), regularizers involving more general differential operators than the gradient have been studied, e.g., in the settings of regularization graphs [9], for higher-order Total Variation operators [12, 10], as well as in structural Total Variation approaches [33] (see also [13] for a review on data-driven approaches in image regularization). Concerning the data-science applicative interest, starting from the variational analysis in [30], a novel research direction has open up for variational studies on point clouds (see, e.g., [6, 15, 16, 24]), as well as related machine-learning applications [14]. The corresponding mathematical study of generalized graph-based differential constraints different from curl $A=0$ is, to the authors’ knowledge, a thriving, mostly unexplored, subject (see, e.g., [34] and the references therein).
The case study we tackle here is a symmetrized counterpart of the nonlocal energies considered in [31, 32]. In order to describe our findings, we first recall the results [31], where the sequence of functionals
@@
In (1.2), the acronym ${\rm GSBV}$ stands for the space of functions with generalised special bounded variation (roughly speaking, maps whose truncations are functions of bounded variations with distributional gradients exhibiting null Cantor part), cf. [4, Section 4.5]. Note that the multiplicative constants in (1.2) depend only on the choice of the integrand $\arctan(x)$ and on the weight function $e^{-|x|^{2}}$.
@@
The structure of (1.3) aligns with the principles of linearized theories in continuum mechanics, where only the symmetric part of the gradient contributes to the deformation energy. This consideration naturally motivates the assumption that only the component of $u(x^{\prime})-u(x)$ projected along the direction $x^{\prime}-x$ should provide energy contributions. By setting $\varepsilon\xi:=x^{\prime}-x$, when the differences $(u(x+\varepsilon\xi)-u(x))\cdot\xi$ are relatively small, the functionals in (1.3) behave like pure bulk energies. For large values of $(u(x+\varepsilon\xi)-u(x))\cdot\xi$, instead, the energies in (1.3) saturate, effectively detecting and penalizing the size of the discontinuities of $u$.
An important remark concerns the fact that the energies in (1.1) only approximate the Mumford-Shah functional under suitable additional assumptions on the geometry of $\Omega$, such as Lipschitz regularity (see [31, Remark 7.1]). Such requirements may not be fully consistent when handling free-boundary problems where discontinuities are already present within the material. Consider, for example, the case of a crack-initiated domain, where the presence of an $(n-1)$-dimensional set $\Gamma\subset\Omega$, representing an initial crack makes the set $\Omega\setminus\Gamma$ not Lipschitz.
@@
In the expression above, $\mathscr{B}$ represents the class of all finite families of pairwise disjoint open balls contained in $\Omega$. This supremum procedure over $\mathscr{B}$ allows the functional $\mathcal{F}^{p}_{\varepsilon}$ to bypass irregularities within the domain $\Omega$. In fact, by choosing $p=1$, the strict superadditivity of the set function $F_{\varepsilon,\xi}(u,\cdot)$ leads to the strict inequality $\mathcal{F}^{1}_{\varepsilon}<\mathcal{F}_{\varepsilon}$.
For $p>1$, the presence of the $L^{p}$-norm in the definition (1.4) should be viewed as a regularizing effect.
The major difficulty we face in the analysis of (1.4) is the lack of a clear compactness strategy. The results of [3, 19, 20] cannot be applied, as the sequence $\mathcal{F}^{p}_{\varepsilon}$ is defined over the space of measurable maps, so that neither a differential structure nor integrability assumptions are given a priori. Thus, our main result is the following compactness and closure theorem for sequences $u_{\varepsilon}$ with equibounded energy $\mathcal{F}^{p}_{\varepsilon}$, for some $p\in[1,+\infty)$.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be open, let $p\geq 1$, and let $\{u_{\varepsilon}\}_{\varepsilon>0}\subset L^{0}(\Omega;\mathbb{R}^{n})$ be such that
@@
$\liminf_{k\to\infty}\mathcal{F}^{p}_{\varepsilon_{k}}(u_{k},\Omega)\geq\mathcal{F}^{p}(u,\Omega):=\int_{\Omega}\varphi_{p}(e(u))\textnormal{d}x+\beta_{p}\mathcal{H}^{n-1}(J_{u}\cup\partial^{*}A),$ (1.7)
@@
where $\varphi_{p}\colon\mathbb{M}^{n\times n}_{sym}\to[0,\infty)$ is a $2$-homogeneous function and $\beta_{p}$ is a positive constant. For $p=1$ we have the following explicit form
$\mathcal{F}^{1}(u,\Omega)=\frac{\pi^{\frac{n}{2}}}{2}\int_{\Omega}\Big{(}|e(u)|^{2}+\frac{1}{2}\textnormal{div}(u)^{2}\Big{)}\,\textnormal{d}x+\frac{\pi^{\frac{n+1}{2}}}{2}\mathcal{H}^{n-1}(J_{u}),$
whenever $u\in\emph{GSBD}(\Omega)$.
In proving Theorem 1.1 we addressed two main correlated hurdles. The first difficulty concerns the identification of a limit map $u$. In this respect, we may compare our setting to the existing literature concerning nonlocal approximation of Free Discontinuity problems. In particular, the recent works [28, 35, 39], inspired by the corresponding results on the Mumford-Shah functional [7, 36], provide a nonlocal approximation of a class of Griffith-type functionals in terms of a bulk energy of the form
featuring the convolution of a volume density $W(e(u))$ for a Sobolev map $u$ and a potential $f$ whose behavior is similar to that of $\arctan$ in (1.3). The compactness issues in the mentioned articles are overcome due to the presence of an a priori differential structure of the displacement $u$, relying on compactness in ${\rm GSBD}$ [19, 20]. We remark that such a direct approach is not applicable in our framework, as $u_{\varepsilon}$ are only measurable fields with no control on their symmetric gradient. A similar issue appeared in a discrete fashion in [1, 23] on a deterministic lattice, which indeed proposed a discrete finite-difference approximation of Griffith-type of energies in the spirit of [5, 17, 18, 38] with a finite range of interaction. Their approach to compactness builds upon the construction of ${\rm GSBD}$-competitors with uniformly bounded energy, which is facilitated by the lack of long-range interactions. This approach allowed again for the application of compactness in ${\rm GSBD}$. The continuous nature of $\mathcal{F}^{p}_{\varepsilon}$ and its infinite horizon make the adaptation of the strategy of [23] rather challenging.
It turned out that in our setting we can work directly on the sequence $u_{\varepsilon}$, bypassing the modification approach of [23]. Our method relies on Fréchet-Kolmogorov, and thus amounts to prove equi-continuity of translations. This technical step is provided in the proof of Theorem 1.1 and shares common ideas with [3], where the proof of the compactness theorem of [19] was first revisited, avoiding the use of Korn and Korn-Poincaré inequalities. Once again, the lack of a symmetric gradient and the structure of $\mathcal{F}_{\varepsilon}$, which features an extra integration over the directions $\xi\in\mathbb{R}^{n}$, do not allow us to select a preferred basis to control translations, as it was the case in [3]. Nevertheless, the slicing properties of the functionals $\mathcal{F}_{\varepsilon}$ make it possible to control the translations of the maps $\arctan(u_{\varepsilon}(x)\cdot\xi)$ both with respect to $x$ and $\xi$. Pointwise convergence out of an exceptional set $A\subseteq\Omega$ is thus recovered by Fréchet-Kolmogorov Theorem.
The next step in the proof of Theorem 1.1 consists in showing that $u$ belongs to ${\rm GSBD}(\Omega)$, yielding the characterization of the domain of the $\Gamma$-limit of $\mathcal{F}^{p}_{\varepsilon}$. Broadly speaking, the non-local nature of the approximating functional $\mathcal{F}^{p}_{\varepsilon}$ results in a limiting function space consisting of measurable vector fields that exhibit generalized bounded deformation in a weaker sense. To explain this phenomenon more precisely, recall that $u\in{\rm GBD}(\Omega)$ if and only if $u$ is a measurable vector field and there exists a finite Radon measure $\lambda$ on $\Omega$ such that, for every $\xi\in\mathbb{S}^{n-1}$, the generalized directional variation $\hat{\mu}_{u}^{\xi}$ of $u\cdot\xi$, as introduced in [26, Definition 4.10], satisfies:
By a slicing argument, a similar approach to that proposed by M. Gobbino in [31] gives that the limit field $u\in L^{0}(\Omega;\mathbb{R}^{n})$ satisfies condition (1.8) in an averaged form with respect to $\xi$, namely,
The fact that (1.8) is replaced in our setting by (1.9) is a huge obstacle to establish the structural properties of the limiting function space. In particular, the technique presented in [26] for deriving the GBD-structural properties fails, and we need to rely on the more recent integral geometric approach introduced in [2]. The main difference from the approach in [2] lies in the role of the flatness of $\mathbb{R}^{n}$, where rescaling and translating the domain preserve straight lines as geodesics, maintaining invariance in the associated GBD-space. Furthermore, the ${\rm GDB}$-theory developed in [25] provides a quite general method to relate the jump set of $u$ to the behavior of codimension one slices, which we are able to adapt to our setting. We remark that such an approach seems not to be successful in a Riemannian manifold. In Propositions 3.143.16, we leverage the above features to avoid relying on a Korn-Poincaré-type inequality to estimate the dimensionality of the set on which the jump of all one-dimensional slices of $u$ concentrate. Ultimately, we conclude in Theorem 3.1 that the limit map $u$ belongs to $\text{GSBD}(\Omega)$, and thus characterize the domain of the $\Gamma$-limit of $\mathcal{F}^{p}_{\varepsilon}$. For the detailed definitions of the relevant quantities and the proof of the structural properties, we refer the reader to Sections 2 and 3, respectively.
An alternative approach. After this paper was submitted, the authors became aware that another equivalent characterization of $GSBD$, based on controlling the measures $\hat{\mu}^{\xi}_{u}$ through finitely many direction $\xi$, has been established in [21]. Instead we adopt here a totally different perspective, by first deriving fine properties of the limiting space and then proving it coincides with $GSBD$ via some integro-geometrical arguments.
Eventually, we obtain the following $\Gamma$-convergence result for the sequence $\mathcal{F}^{p}_{\varepsilon}$.
@@
Let $\Omega\subset\mathbb{R}^{n}$ be open. For every $p\geq 1$ the family of functionals $\{\mathcal{F}^{p}_{\varepsilon}\}_{\varepsilon>0}$ $\Gamma$-converges in $L^{0}(\Omega;\mathbb{R}^{n})$ to the Griffith-type functional $\mathcal{F}^{p}$ defined as
@@
Combining Theorem 1.1 with the $\Gamma$-convergence of Theorem 1.2, we further deduce convergence of minimisers of $\mathcal{F}_{\varepsilon}^{p}$ to minimisers of $\mathcal{F}^{p}$ under suitable Dirichlet boundary conditions (cf. Theorem 5.2).
@@
Organization of the paper. The paper is organized as follows. In Section 2, we collect a few preliminary definitions and results. Section 3 is devoted to establish the main properties of the spaces of limiting deformations. Sections 4 and 5 are devoted to the proofs of our compactness, and $\Gamma$-convergence result, respectively.
@@
Outlook. The analysis provided in this work offers the opportunity to explore several further research lines. On the one hand, the compactness argument of Theorem 1.1 and the slicing technique of Theorem 1.2 may be further tested against discrete finite-difference approximations on point clouds of vectorial free discontinuity problems, in the spirit of [8, 16, 30] for the Total Variation, the Mumford-Shah, and the perimeter functionals. Such research line is the subject of a preprint which is currently in preparation.
On the other hand, the use of a slicing argument for the study of the variational limit restricts the choice of the approximating sequence $\mathcal{F}_{\varepsilon}$, which in turn determines the class of admissible densities $\varphi_{p}$ in (1.10). The study of more general approximations and of integral representation formulas in the spirit of [22, 1, 23] will be the subject of future investigation. A possible strategy to generalize our result to the case of arbitrary Lamé constants could be to rely on the presence of further discrete-divergence terms, along the lines of [1, 23]. The extension of our compactness result and of our identification of the limiting domain for such a model would be possible without any change to the structure of the current proof. The associated Gamma-convergence analysis, on the other hand, deserves the development of a novel representation theory.